OpenStudy (anonymous):

Solve the equation. Show your work. 16^3 = 4^x

OpenStudy (anonymous):

16^3 = 4^x ln 16^3 = ln 4^x ln 16^3 = x ln4 x=(ln 16^3)/(ln4) = (ln 4^6)/(ln 4) = (6ln4)/(ln4) =6